# T-MAE: Temporal Masked Autoencoders for Point Cloud Representation Learning

### v1

@@ 4.3 Ablation study
In this section, we perform a series of ablation studies to assess the necessity of various components and justify the selection of hyperparameters. To cost-effectively experiment, the split 0 with 5% data of the data-efficient benchmark is used to finetune the model by default.
Is a delicate fusion module necessary? To compare with a model taking two frames as input, we modified the input of GD-MAE [52] by concatenating all points of two consecutive frames. We pre-trained, finetuned, and evaluated the modified GD-MAE with the same setting as T-MAE. As depicted in Table 2, directly merging two frames improves the metrics in terms of vehicles and pedestrians, which probably results from the density of target objects being doubled, as shown in Fig. 3. However, it has a negative impact on cyclists, resulting in a decrease in overall performance. The drop for cyclists might be attributed to the drift of estimated bounding boxes caused by the fast velocity of bicycles and their relatively small dimensions. In contrast, T-MAE significantly improves the performance thanks to the delicate pre-training strategy and fusion mechanism.
@@
Table 3: Ablation study on model architecture. The proposed SiamWCA demonstrates superior performance in terms of overall mAP and mAPH.
Encoder design and fusion mechanism. The proposed architecture consists of a Siamese encoder and a WCA module, which is denoted as (e) in this ablation study. In addition, several variants are compared: (a) Asymmetric encoder: The encoder for the previous frame is replaced with a more lightweight encoder, resulting in a reduction of the number of channels by half. Consequently, the two encoders no longer share weights. (b) We employ the SimSiam [7] approach, where one of the two encoders is disconnected from the computational graph, preventing it from being updated by gradient propagation. The encoder for the current frame is updated by gradients and then shares its weights with the detached encoder. (c) The design of the Siamese encoder stays unaltered, while a typical cross-self attention module is adopted in a window-based manner. These three variants are depicted in Fig. 5. (d) Instead of using GD-MAE [52] as our single-frame backbone, the standard (ST) MAE [18] is adopted to 3D space and serves as the Siamese encoder. The superiority of the proposed SiamWCA is shown in Tab. 3. Variant (c) outperforms SiamWCA in terms of pedestrians but at the expense of cyclists’ performance.
(a) Asym. Enc. (b) SimSiam Enc. (c) WCA + WSA
@@
Figure 5: Three architecture variants of our SiamWCA backbone. (a) Asymmetric encoders: the encoder for the previous frame is replaced by a more lightweight encoder. (b) SimSiam-style [7] encoder: one encoder receives no gradient updates. (c) An additional windows-based self-attention is attached as the classic Transformer [43].
Previous frame L2 Overall Vehicle Pedestrian Cyclist
mAP mAPH AP APH AP APH AP APH
$[-5,-1]$ 49.86 47.61 56.70 56.16 59.47 54.65 33.41 32.03
$-1$ 49.45 46.78 56.56 56.02 57.96 51.94 33.82 32.37
$-3$ 50.20 48.07 56.71 56.18 60.19 55.71 33.69 32.33
$-5$ 49.87 47.69 56.73 56.20 59.91 55.26 32.98 31.61
Table 4: Temporal gap for inference. Different temporal gaps were tested for inference. $[-5,-1]$ indicates a random frame is sampled from the last five frames. Overall, a three-frame temporal gap worked best in our experiments.
Temporal interval for inference. This ablation study explores the impact of different temporal gaps during inference on model performance. We establish a baseline with a random interval between two frames. Specifically, for each frame, the previous frame is selected randomly from the preceding five frames. Furthermore, the temporal gap is fixed to $-1$, $-3$, and $-5$, respectively. If the number of available frames is less than the predefined threshold, the earliest available frame is selected. Tab. 4 illustrates that optimal performance is achieved with a fixed temporal gap of three. Specifically, the model excels in recognizing vehicles, pedestrians, and cyclists when the temporal gap is set to $-5$, $-3$, and $-1$, respectively. This observation highlights the varied impact of different temporal gaps on distinct object categories. Notably, while the performance remains relatively stable for vehicles and cyclists, there is a pronounced improvement in pedestrian recognition as the temporal gap transitions from $-1$ to $-3$. This highlights the considerable advantage of including more distant historical frames to achieve accurate pedestrian detection as well as his/her heading.
@@ 5 Conclusion
In this paper, we introduced the Temporal Mask AutoEncoders (T-MAE), a novel self-supervised paradigm for LiDAR point cloud pre-training. Building upon the single-frame MAE baseline, we incorporated historical frames into the representation using SiamWCA, with the proposed WCA module playing a pivotal role. This pre-training enabled the model to acquire robust representations and the ability to capture motion even with very limited labeled data. By considering temporally distant frames instead of consecutive ones, we achieved additional performance improvements. Our experiments on the Waymo Open dataset demonstrate the effectiveness of our approach by showing mostly improvements over state-of-the-art methods.
6 Acknowledgements
@@
This work was financially supported by TomTom, the University of Amsterdam and the allowance of Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy. Fatemeh Karimi Nejadasl was financed by the University of Amsterdam Data Science Centre.
Data Initialization Overall Vehicle Pedestrian Cyclist
amount L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
5% Random 49.59 46.15 54.42 53.87 52.59 44.17 41.76 40.41
@@
T-MAE (Ours) 59.18 ${}^{+1.74}$ 57.29 ${}^{+2.81}$ 59.66 59.15 63.90 59.87 53.97 52.84
Table 5: Performance comparison on the Waymo validation set [39]. Results for other methods are taken from MV-JAR [48]. All methods are finetuned with Subset 1.
Data Initialization Overall Vehicle Pedestrian Cyclist
amount L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
5% Random 43.33 40.10 55.39 54.81 55.75 47.69 18.86 17.79
@@
T-MAE (Ours) 58.15 ${}^{+1.34}$ 56.26 ${}^{+2.29}$ 60.08 59.56 66.02 61.82 48.34 47.41
Table 6: Performance comparison on the Waymo validation set [39]. Results for other methods are taken from MV-JAR [48]. All methods are finetuned with Subset 2.
Data Initialization Overall Vehicle Pedestrian Cyclist
amount L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH L2 mAP L2 mAPH
5% Random 45.53 42.18 54.62 54.06 53.93 45.54 28.05 26.94
@@
T-MAE (Ours) 58.82 ${}^{+2.06}$ 56.87 ${}^{+3.01}$ 60.01 59.50 64.51 60.31 51.94 50.79
Table 7: Average performance on the Waymo validation set [39], averaged across the models finetuned with Subset 0--2. Results for other methods are taken from MV-JAR [48]. T-MAE consistently enhances model performance with a notable margin compared to random initialization. Furthermore, T-MAE attains SOTA average performance across all metrics when finetuned with 10% data.
(a) Input (b) Attention scores for prior scan
@@
Figure 6: Visualization of attention scores. (a) The input consists of two point clouds: the entire previous scan (red points) and the current scan (green points) that contains only points within the ground-truth bounding boxes. Note that, the blue bounding boxes in (a) serve solely for visualization purposes and do not function as input. (b) The pillar-wise attention scores are visualized. The attention scores are derived from the WCA module and mapped to a colormap ranging from black to white. The primary attention is placed on the target objects from the previous scan. This implies that the WCA is able to locate corresponding objects. Notably, WCA is capable of accurately trace to the source object, which is manually indicated by a skyblue box, even in cases where the vehicle is moving.
@@
Random/T-MAE 5% 30 7.9 2.37 4.0 40.29/48.07 44.76/55.71
10% 30 15.8 4.75 7.9 53.13/57.05 53.04/59.24
20% 30 31.6 9.50 15.8 57.61/61.72 58.41/64.53
Table 8: Quantitative details on comparison of finetuning iterations. Our scheme requires much less iterations for finetuning compared to MV-JAR [48].
Appendix A Additional Results on Waymo Subsets
The data-efficient benchmark, proposed by MV-JAR [48], contains three splits, each of which contains four tiers of data, namely 5%, 10%, 20% and 50%. In the context of data organization, each split corresponds to a subdivision where a lower level of data is a subset of a higher level. There is no overlap observed between different splits for the data corresponding to the 5% and 10% subsets. In the main paper, the majority of experiments were performed by finetuning the model with the split 0. To alleviate the variance caused by being finetuned with a small amount of data, we further present the performance obtained when the model is finetuned using the remaining two splits, as shown in Tab. 5 and Tab. 6. Furthermore, the average results are presented in Tab. 7. In conclusion, it can be observed that the model performance exhibits variation when finetuned with different subsets. However, T-MAE consistently demonstrates a noteworthy improvement in model performance compared to random initialization. Additionally, T-MAE attains SOTA average performance across all metrics when finetuned using 10% data.
Appendix B Visualization of cross-attention scores
To understand the knowledge acquired by the SiamWCA during pre-training, we employ attention visualization to identify critical regions of the previous scan. In particular, the entire previous scan and the target objects of current scan are input into the SiamWCA. More precisely, the input to the model for the current scan solely consists of points contained within the 3D ground-truth bounding boxes. The purpose of removing other points is to identify the specific areas of emphasis within the WCA module when the queries are solely target objects, e.g. vehicles and pedestrians. As illustrated in Fig. 6 (b), the attention is mainly attached to target objects, indicating that the WCA successfully detects and localizes these entities in the previous frame. Furthermore, the presence of vehicular action is perceptible even if the WCA module is only trained with unlabeled data, indicating the effectiveness of our T-MAE pre-training strategy.
Appendix C Comparison of training iterations.
MV-JAR [48] applies varying numbers of epochs during finetuning, depending on the data amount. Rather than varying numbers, we employ a predetermined number of epochs, leading to significantly less iterations for finetuning without performance drop. Detailed information is presented in Tab. 8. The total number of iterations is calculated by multiplying the number of input scans by the number of epochs. As depicted in Fig. 1 and Tab. 8, T-MAE achieve better performance than MV-JAR [48] while requiring a smaller number of finetuning iterations. In addition, our scheme requires just one GT annotation every two input frames. As a result, the total number of iterations would be even smaller if it is calculated by multiplying the number of GT by the number of epochs.

### v3

@@ 4.3 Ablation Study
We perform ablation studies on the Waymo dataset [49] to justify design choices and hyperparameters. For cost-effective experiments, the split 0 with 5% data of the data-efficient benchmark [64] is used to finetune the model by default.
Is a delicate fusion module necessary?
@@
To compare with a model taking two frames as input, we modified the input of GD-MAE [64] by concatenating all points of two consecutive frames. We pre-trained, finetuned, and evaluated the modified GD-MAE with the same setting as T-MAE. As depicted in Table 3, directly merging two frames improves the metrics in terms of vehicles and pedestrians, which probably results from the density of target objects being doubled, as shown in Fig. 2. However, it has a negative impact on cyclists, resulting in a decrease in overall performance. The drop for cyclists might be attributed to the drift of estimated bounding boxes caused by the fast velocity of bicycles and their relatively small dimensions. In contrast, integrating a historical frame by our method consistently improves overall and class-specific metrics compared to the single-frame baseline.
Table 4: Ablation study on model architecture. The proposed SiamWCA (e) demonstrates superior performance in terms of overall mAP and mAPH.
@@
(a) Asym. Enc. (b) SimSiam Enc. (c) WCA + WSA (d) WSA
@@
Figure 6: Four architecture variants of our SiamWCA backbone. (a) Asymmetric encoders: the encoder for the previous frame is scaled down. (b) SimSiam-style [10] encoder: one encoder receives no gradient updates. (c) An additional windows-based self-attention is attached as the classic Transformer [55]. (d) The fusion is implemented by a concat-WSA-split operation.
Backbone design. The proposed architecture consists of a Siamese encoder and a WCA module. We ablate these components with alternatives. (a) Asymmetric encoder: it derives from a Siamese encoder but the encoder for $\mathcal{P}^{1}$ is scaled down. Consequently, the two encoders no longer share weights. (b) We employ the SimSiam [10] approach, where one of the two encoders is detached from the computational graph, preventing it from being updated by gradient propagation. The encoder for the current frame is updated by gradients and then shares its weights with the detached encoder. (c) The design of the Siamese encoder stays unaltered, while a typical cross-self attention module is adopted in a window-based manner. (d) The WCA module in our SiamWCA is replaced with a windowed self-attention (WSA) module. However, since self-attention only needs one input, the tokens of both frames have to be concatenated. Then, the enhanced tokens of the current frame are split from the output of WSA module. These variants are depicted in Fig. 6. The superiority of the proposed SiamWCA is shown in Tab. 4. Variant (c) outperforms SiamWCA in terms of pedestrians but at the expense of both cyclists’ performance and extra parameters. Therefore, based on the comprehensive comparison, a Siamese encoder and a WCA fusion module are selected.
Compatibility. To verify that the proposed SSL method T-MAE is independent of a specific encoder or detector, we replace the encoder SPT [64] with two other encoders, i.e. SST [16] and SpCNN [19], and the detector with a two-stage detector, Graph R-CNN [65]. Table 5 shows that T-MAE constantly improves performance by a significant margin, showcasing its compatibility.
Table 5: Ablation study on the encoder and the detector. The original encoder SPT [64] is replaced with SST [16] and SpCNN [19] and the original detector enterPoint [68] is replaced with Graph R-CNN [65]. Due to space limits, only overall performance is presented.
Data Amount Encoder Detector Random Init. T-MAE (Ours)
mAP mAPH mAP mAPH
5% SPT CenterPoint 43.68 40.29 51.47 49.46
SST CenterPoint 44.26 41.19 51.59 49.24
SpCNN CenterPoint 46.02 43.31 52.08 49.95
SPT Graph R-CNN 51.18 47.27 56.92 54.70
100% SPT CenterPoint 71.30 69.13 72.30 70.52
SPT Graph R-CNN 72.76 70.04 75.16 73.50
Temporal interval for inference. The temporal interval between two frames should be constrained, as previously discussed in Sec. 3.2. To investigate this matter, we conducted experiments on the Waymo Dataset [49], suggesting that a fixed interval of 0.3 seconds is better than using two consecutive frames. Additional information is available in Appendix Sec. S3.
@@ 5 Conclusion
We introduced Temporal Masked Autoencoders (T-MAE), a novel self-supervised paradigm for LiDAR point cloud pre-training. Building upon the single-frame MAE baseline, we incorporated historical frames into the representation using SiamWCA, with the proposed WCA module playing a pivotal role. This pre-training enabled the model to acquire robust representations and the ability to capture motion even with very limited labeled data. By constraining the temporal interval of two frames, we achieved additional performance improvement. Our experiments on the Waymo dataset and the ONCE dataset demonstrate the effectiveness of our approach by showing improvements over state-of-the-art methods.
Acknowledgements
@@
This work was financially supported by TomTom, the University of Amsterdam and the allowance of Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy. This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-7940. Fatemeh Karimi Nejadasl was financed by the University of Amsterdam Data Science Centre.
S1 Implementation Details
We follow the training settings of GD-MAE [64]. Some important configurations are listed in Tab. S1 and Tab. S2. We use the same masking ratio, the per-pillar number of predicted points $K^{O}$ and target reconstructed points $K^{GT}$ as GD-MAE because these parameters have negligible effects according to the ablation studies in GD-MAE [64].
Table S1: Training details.
Config Pre-training Finetuning
optimizer AdamW [36]
optimizer momentum $\beta_{1},\beta_{2}=0.9,0.99$
weight decay 0.01
max learning rate 0.003
learning rate scheduler a cyclic learning rate
with cosine annealing
batch size (Waymo [49]) 4 3
batch size (ONCE [37]) 8 6
epoch 12 30
masking ratio 0.75 -
# predicted points $K^{O}$ 16 -
# target reconstructed points $K^{GT}$ 64 -
Table S2: Dataset-specific details. Finetuning time indicates the duration of finetuning using the entire training set.
Config Waymo [49] ONCE [37]
window size $(8,8,1)$
detection range - x-axis (m) $(-74.88,74.88)$
detection range - y-axis (m) $(-74.88,74.88)$
detection range - z-axis (m) $(-2,4)$ $(-3,5)$
pillar size (m) $(0.32,0.32,6)$ $(0.32,0.32,8)$
temporal batch 6 3
GPUs 8 $\times$ Tesla V100 4 $\times$ A100 (40GB)
Pre-training time (GPU day) 8 $\times$ 4 4 $\times$ 4.5
Finetuning time (GPU day) 8 $\times$ 5 4 $\times$ 0.25
Table S3: Ablation experiments on the Waymo dataset [49].
Ablation Target Setting L2 Overall
mAP mAPH
Two-frame alignment baseline 41.42 37.56
w/ alignment 44.05↑2.63 41.28↑3.72
Data augmentation w/ copy-n-paste [63] 46.76↑5.34 43.93↑6.37
Length of temporal batch 6$\rightarrow$3 45.88↑4.46 41.96↑4.40
6 46.76↑5.34 43.93↑6.37
6$\rightarrow$12 44.96↑3.54 41.92↑4.36
S2 Additional Ablation Study
Additional results on the Waymo dataset [49] are provided to justify design choices and hyperparameter values. Note that, in this section, we use 20% of the training set to pre-train our model and use 5% of the training set to fine-tune it for cost-effective experiments.
Two-frame Alignment. We start with using SiamWCA as a baseline and investigate whether it is necessary to align the previous frame to the coordinate system of the current frame. As shown in Tab. S3, aligning two frames significantly improves the metrics.
Data Augmentation. Based on the baseline with two-frame alignment, we further investigate the effect of a data augmentation technique, namely copy-n-paste [63]. Specifically, it first generates a ground truth instance database during dataset pre-processing. Then, during finetuning, several ground truth instances are randomly placed into the scene. This augmentation boosts the detection rates. Note that only ground truths in the 5% split are included in the database, which avoids ground truth leakage.
Length of Temporal Batch. This hyperparameter is set as 6 for the previous experiments in Tab. S3. In this ablation, we pre-train and fine-tune the model with different lengths of the temporal batch. As shown in Tab. S3, setting the length of the temporal batch as 6 achieves the optimal performance.
S3 Temporal interval for inference.
Figure S1: Temporal interval for inference. Different temporal intervals were tested for inference. $[0.1,0.5]$ indicates a random interval between $0.1$ and $0.5$ seconds. A fixed interval ranging from 0.1 to 0.9 seconds was also tested. Overall, a fixed interval of 0.3 seconds works best in our experiments.
This ablation study explores the impact of different temporal intervals on the model performance during inference. We establish a baseline with a random interval ranging from 0.1 to 0.5 seconds. In other words, $\mathcal{P}^{t_{1}}$ is selected randomly from the last five frame of $\mathcal{P}^{t_{2}}$. As a comparison, the temporal interval is fixed to values ranging from $0.1$ to $0.9$ seconds. If there is no satisfactory frame available, the earliest available frame is selected as $\mathcal{P}^{t_{1}}$. As shown in Fig. S1, The AP and APH for pedestrians achieve relatively high values when the interval is 0.3 and 0.5 seconds, while those for cyclists hit optimal when the interval is 0.1 seconds. The different optimal intervals could be attributed to the different speeds of the two categories. Pedestrians can barely move within 0.1 seconds, leading to historical information less useful. On the contrary, cyclists move relatively faster and thus the same bicycle of two frames is easy to be recognized as two instances if the temporal gap is large. Fig. S1 also illustrates that optimal overall performance is achieved when the interval is 0.3 seconds. Therefore, the previous third frame, which corresponds to a temporal interval of 0.3 seconds, is selected as $\mathcal{P}^{t_{1}}$ for the Waymo dataset.
S4 Additional Results on Waymo dataset
Table S4: Performance comparison on the Waymo validation set [49]. All methods are fine-tuned with 20% uniformly sampled frames, namely sampling one frame per five frames. Random initialization denotes training from scratch. Differences between T-MAE pre-training and random initialization are highlighted in red. ∗∗ indicates results from [64]. Other results are from AD-PT [70] or the survey [17]. Best results are highlighted as first, second, and third.
Data Amount Initialization Overall Vehicle Pedestrian Cyclist
mAP mAPH mAP mAPH mAP mAPH mAP mAPH
ProposalContrast [67] 66.67 64.20 65.22 64.80 66.40 60.49 68.48 67.38
BEV-MAE [33] 66.70 64.25 64.72 64.22 66.20 60.59 69.11 67.93
Occupancy-MAE [39] 65.86 63.23 64.05 63.53 65.78 59.62 67.76 66.53
MAELi-MAE [29] 65.60 63.00 64.22 63.70 65.93 59.79 66.66 65.52
GD-MAE [64]∗∗ 70.24 67.14 67.67 67.22 73.18 65.50 69.87 68.71
AD-PT [70] 67.17 64.65 65.33 64.83 67.16 61.20 69.39 68.25
T-MAE (Ours) 70.92↑1.89 69.11↑2.63 68.07 67.61 74.38 70.56 70.32 69.15
Table S5: Performance on the Waymo validation set [49]. Results for other methods are taken from MV-JAR [60]. All methods are finetuned with Subset 1.
Finetuning Initialization Overall Vehicle Pedestrian Cyclist
split mAP mAPH mAP mAPH mAP mAPH mAP mAPH
5% S1 Random 49.59 46.15 54.42 53.87 52.59 44.17 41.76 40.41
@@
T-MAE (Ours) 59.24↑1.80 57.26↑2.78 59.71 59.19 64.44 60.29 53.58 52.31
Table S6: Performance on the Waymo validation set [49]. Results for other methods are taken from MV-JAR [60]. All methods are finetuned with Subset 2.
Finetuning Initialization Overall Vehicle Pedestrian Cyclist
split mAP mAPH mAP mAPH mAP mAPH mAP mAPH
5% S2 Random 44.25 41.48 53.04 52.53 55.16 49.42 24.56 22.48
@@
T-MAE (Ours) 58.64↑1.83 56.71↑2.73 60.20 59.70 66.10 61.79 49.62 48.63
Table S7: Average performance on the Waymo validation set [49], averaged across the models finetuned with Subset 0$\sim$2. Results for other methods are taken from MV-JAR [60]. T-MAE consistently enhances model performance with a notable margin compared to random initialization.
Finetuing Initialization Overall Vehicle Pedestrian Cyclist
split mAP mAPH mAP mAPH mAP mAPH mAP mAPH
[…]
@@
T-MAE (Ours) 59.27↑2.50 57.32↑3.46 60.06 59.55 65.26 61.06 52.50 51.34
We explore two methods to obtain a subset of the training set for finetuning, but all models are evaluated on the same validation set. Uniform Sampling is commonly used [67, 64, 32]. Specifically, frames from all sequences are concatenated, from which the target frames are sampled with a fixed interval. Data-efficient Benchmark [60] selects a certain number of sequences as a training subset instead of uniform sampling frames, which aims to solve the data diversity issue existing in uniform sampling. More specifically, while finetuned on uniformly sampled frames, models always converge to a similar performance if they are finetuned for adequate iterations no matter how many percentages of labelled data are used. As a result, fewer data do not lead to a shorter finetuning time. The data-efficient benchmark enables models to converge with much fewer iterations. Therefore, to evaluate pre-trained representation more efficiently, we conduct experiments on the data-efficient benchmark, as presented in Tab. 1.
Results Analysis. Under the uniform sampling setting, we finetuned the T-MAE pre-trained model with 20% uniformly sampled frames. As shown in Tab. S4, T-MAE outperforms other methods, which aligns with the statement made in the main paper. Under the data-efficient setting, the 5% and 10% splits contain limited samples, which may lead to performance variance. Therefore, except for the split 0 used for Tab. 1, we also compare models in Tab. S5 and Tab. S6 when they are finetuned with split 1 and split 2, both of which are provided by MV-JAR [60] as well. Furthermore, the average results are presented in Tab. S7. In conclusion, it can be observed that the model performance exhibits variation when finetuned with different subsets. However, T-MAE pre-training consistently improves model performance compared to random initialization. T-MAE pre-training surpasses other SSL methods in terms of most class-specific metrics and all overall metrics. Notably, T-MAE outperforms other methods in terms of all metrics while the results are averaged (see Tab. S7).
S5 Transferring performance.
To assess the transferability of T-MAE pre-training, we pre-train the model on the ONCE dataset [37] and then fine-tune it on the Waymo dataset [49]. Table S8 demonstrates that T-MAE offers robust generalizability and transferability.
Table S8: Transferability comparison.
Initialization Data Overall Vehicle Pedestrian Cyclist
Random - 66.48 66.15 66.76 66.54
GD-MAE [64] ONCE 66.61 67.18 64.82 67.83
T-MAE (Ours) ONCE 67.86 68.06 67.25 68.26
S6 Multi-frame comparison with non-SSL methods.
Since there are no existing multi-frame SSL methods, we compare our approach with robust non-SSL baselines that utilize multi-frame as inputs. Table S9 demonstrates our method surpasses other supervised methods while requiring only two frames, highlighting the effectiveness of our method.
Table S9: Multi-frame comparison with non-SSL methods. “-” indicates not available.
Method Frames L2 Overall Vehicle Pedestrian Cyclist
mAP mAPH mAP mAPH mAP mAPH mAP mAPH
3D-MAN [66] 16 - - 67.6 67.1 62.6 59.0 - -
CenterPoint [68] 4 70.8 69.4 69.1 68.6 71.7 68.6 71.6 70.9
SST [16] 3 - - 68.5 68.1 75.1 70.9 - -
T-MAE (Ours) 2 72.3 70.5 69.4 68.9 75.8 72.0 71.8 70.7
(a) Input (b) Attention scores for prior frame
@@
Figure S2: Visualization of attention scores. (a) The input consists of two point clouds: the entire previous frame (red points) and the current frame (green points) that contains only points within the ground-truth bounding boxes. Note that, the blue bounding boxes in (a) serve solely for visualization purposes and do not function as input. (b) The pillar-wise attention scores are visualized. The attention scores are derived from the WCA module and mapped to a colormap ranging from black to white. The primary attention is placed on the target objects from the previous frame. This implies that the WCA is able to locate corresponding objects. Notably, WCA is capable of accurately trace to the source object, which is manually indicated by a skyblue box, even in cases where the vehicle is moving.
S7 Attention Learned by T-MAE Pre-training
In the main paper, T-MAE pre-trained weights are loaded to both the Siamese encoder and WCA module. We conduct an ablation study where only the Siamese encoder is initialized by the pre-trained weights and the WCA module is randomly intialized. As shown in Tab. S10, initializing WCA with T-MAE pre-trained weights significantly improves mAPH for pedestrians, which indicates a way better direction detection for pedestrians. It also boosts metrics for cyclists with a big margin.
Table S10: Ablation on WCA initialization. SE stands for the Siamese Encoder.
Initialization Pre-trained Overall Vehicle Pedestrian Cyclist
SE WCA mAP mAPH mAP mAPH mAP mAPH mAP mAPH
Random 43.68 40.29 54.05 53.50 53.45 44.76 23.54 22.61
Partially random 48.19↑4.51 45.16↑4.87 55.91 55.38 56.29 48.74 32.38 31.36
T-MAE (ours) 51.47↑7.79 49.46↑9.17 57.13 56.63 59.69 55.28 37.61 36.48
To further understand the knowledge acquired by the WCA module during pre-training, we employ attention visualization to identify critical regions of the previous frame. In particular, the entire previous frame and the target objects of the current frame are input into the SiamWCA that loads T-MAE pre-trained weights. More precisely, the input to the model for the current frame solely consists of points contained within the 3D ground-truth bounding boxes. The purpose of removing other points is to identify the specific areas of emphasis within the WCA module when the queries are solely target objects, e.g. vehicles, pedestrians, and cyclists. As illustrated in Fig. S2 (b), the attention is mainly attached to target objects, indicating that the WCA successfully detects and localizes these entities in the previous frame. Furthermore, the presence of vehicular action is perceptible even if the WCA module is only trained with unlabeled data, indicating the effectiveness of our T-MAE pre-training strategy.
S8 Comparison of training iterations
Table S11: Quantitative details on comparison of finetuning iterations. Our scheme requires much less iterations for finetuning compared to MV-JAR [60].
@@
Random/T-MAE 5% 30 7.9 2.37 4.0 40.29/49.46 44.76/55.28
10% 30 15.8 4.75 7.9 53.13/57.99 53.04/61.10
20% 30 31.6 9.50 15.8 57.61/61.80 58.41/64.66
MV-JAR [60] applies varying numbers of epochs during finetuning, depending on the data amount. Rather than varying numbers, we employ a predetermined number of epochs, leading to significantly fewer iterations for finetuning without a performance drop. Detailed information is presented in Tab. S11. The total number of iterations is calculated by multiplying the number of input scans by the number of epochs. As depicted in Fig. 1 and Tab. S11, T-MAE outperforms MV-JAR [60] while requiring a smaller number of finetuning iterations.
S9 Qualitative Results
Fig. S3 shows the qualitative results of our method on the Waymo dataset [49]. Fig. S4 and Fig. S5 illustrates qualitative results on the ONCE datset [37]. Our method generates accurate bounding boxes even if the scene is complex.
S10 Limitations
While our work has achieved encouraging results, there is space for further improvements. For instance, the alignment of two frames relies on their transformation matrices; without proper alignment, there is a drop in performance. If this could be eliminated, the proposal would be more efficient. Another shortcoming is that the transformer blocks are computationally heavy, which restricts the window size. This leads to a smaller receptive field and thus makes the network easy to lose track of fast-moving objects. The transformer-based encoder also increases the inference time, which was tested on a single NVIDIA RTX 3090 and measured as 429 ms per frame. Consequently, the network is not real-time.
Figure S3: Qualitative results on the Waymo dataset [49]. We depict ground truth and predictions as boxes colored in red and green for several exemplary scenes.
Figure S4: Qualitative results on the ONCE dataset [37]. We depict ground truth and predictions as boxes colored in red and green for several exemplary scenes.
Figure S5: Qualitative results on the ONCE dataset [37]. We depict ground truth and predictions as boxes colored in red and green for several exemplary scenes.